### v1

@@ 2.3 Sampling Algorithm for SENT
Thm. 2.11 offers a simple strategy to sample a causal and Hamiltonian SENT: one needs to traverse the graph and choose the neighborhood set as all neighbors of the current node that have been visited. The traversing strategy could be achieved through a random path sampling or a depth-first search. In Algorithm 1, we provide a sampling strategy based on random path sampling with breaks.
@@
4:$s\leftarrow[]$
5:$v\leftarrow\texttt{RandomSample}(U)$; $U\leftarrow U\setminus\{v\}$
6:$t\leftarrow[(v,\emptyset)]$triangleright first neighborhood trail
@@
Complexity analysis.
The length of a SENT including the sizes of neighborhood sets (in other words, tokenized SENT defined in Section 2.4) is bounded by the number of edges as it can visit each edge only once. Therefore, both the time and space complexity of sampling a SENT from graph $G$ are $\mathcal{O}(m)$ where $m$ is the number of edges.
2.4 Tokenization of SENT
Previous works have explored related concepts of sequences in graphs. For example, You et al. (2018) investigated causal Hamiltonian neighborhood sequences generated through breadth-first search, while Liao et al. (2019); Goyal et al. (2020) constructed SENT-like sequences using depth-first search. However, neither of these works interpreted these sequences as a language. Here, we present a method to bridge the gap between graph generation and language modeling.
The tokenization process starts by mapping all isomorphic SENTs to the same sequence, by reindexing the vertices according to their first occurrence order within the sequence. Specifically, if we denote this ordering function for a SENT $s$ by $\pi:V_{s}\to\{1,\dots,|V_{s}|\}$, $s$ is then replaced with its ordered representation $\pi(s)$. Thanks to the isomorphism property of SENT (Thm. 2.4), $\pi(s)$ generates a graph isomorphic to $G_{s}$ while ensuring the obtained sequence invariant to the node ordering of the input graph.
To convert an (ordered) SENT into a machine-readable sequence, we tokenize it into a sequence of indices using special tokens. These tokens include symbols such as ‘ / ’ to indicate a breakage between segments, and ‘$\bm{<}$’ and ‘$\bm{>}$’ to mark the start and end of a neighborhood set. Specifically, for any $s:=(s_{1},\dots,s_{k})\in\mathcal{S}^{\mathcal{N}}$, we define the tokenization function Token as follows:
$\texttt{Token}(s):=\texttt{Token}(s_{1})\mathbin{\|}[~{}\textbf{/}~{}]\mathbin{\|}\cdots\mathbin{\|}[~{}\textbf{/}~{}]\mathbin{\|}\texttt{Token}(s_{k}),$
where
$\texttt{Token}(s_{i}):=\mathbin{\|}_{w\in s_{i}}\texttt{Token}(w),$
@@ 2.5 Extension to Attributed Graphs
Our method can be easily extended to graphs with categorical (or discretized) attributes by inserting node and edge attributes in an interleave fashion into the tokenized SENT sequence. Specifically, let $L_{\mathrm{node}}(v)$ and $L_{\mathrm{edge}}(u,v)$ be the attributes of a node $v$ and an edge $(u,v)$ respectively. Using the same notation as above, we define for any $s_{i}:=(w_{1},\dots,w_{q})\in{\mathcal{T}^{\mathcal{N}}}$ with $w_{i}=(v_{i},\cdot)$:
$\texttt{Token}(s_{i})$ $:=\texttt{Token}(w_{1})\mathbin{\|}[L_{\mathrm{edge}}(v_{1},v_{2})]\mathbin{\|}\texttt{Token}(w_{2})$
$\mathbin{\|}\dots\mathbin{\|}\texttt{Token}(w_{q}),$
$\texttt{Token}(w)$ $:=[v,L_{\mathrm{node}}(v),\bm{<},L_{\mathrm{edge}}(v,u_{1}),u_{1},$
$\dots,L_{\mathrm{edge}}(v,u_{p}),u_{p},\bm{>}].$
@@ 2.6 Autoregressive Modeling of Tokenized SENTs
where the conditional probability $p_{\theta}$ is modeled using a neural network with parameters $\theta$. The architecture of the neural network can be any SOTA sequence model.
@@ 3 Related Work
Autoregressive models generate graphs by sequentially adding nodes and edges. GraphRNN (You et al., 2018) pioneered this approach by framing graph generation as a sequence prediction task, demonstrating the capacity of recurrent neural networks (RNNs) (Chung et al., 2014) to capture complex structures. DeepGMG (Li et al., 2018) introduced a probabilistic policy framework for conditional generation, while GRAN (Liao et al., 2019) and BiGG (Dai et al., 2020) enhanced efficiency and scalability by generating multiple nodes and edges in parallel.
Recent research has focused on optimizing the generation order. Chen et al. (2021) highlighted that the ordering of node and edge additions impacts graph quality, and GraphARM (Kong et al., 2023) applied reinforcement learning to dynamically refine this order. Goyal et al. (2020) incorporated logical constraints to improve domain-specific generation, and Bacciu et al. (2020) proposed Bayesian reasoning to better capture graph dependencies.
Although these models have shown to be efficient and effective in synthetic datasets, they face inherent unification limitations, reducing their applicability to large-scale real-world datasets. Our proposed unified and powerful sequence representation of graphs aims to address these challenges.
@@
Other graph generative models include variational, GAN-based, and diffusion-based approaches. GraphVAEs (Kipf & Welling, 2016; Simonovsky & Komodakis, 2018) employ variational autoencoders to learn continuous latent representations, effectively generating small graphs but struggling with more complex structures. GAN-based models, such as NetGAN (Bojchevski et al., 2018) and SPECTRE (Martinkus et al., 2022), generate graphs by modeling graph descriptors like random walks and spectral features.
Diffusion-based models iteratively refine noise into structured graphs through reverse diffusion steps. Continuous diffusion models (Niu et al., 2020; Jo et al., 2022) adapt denoising diffusion probabilistic models for graph generation. To leverage graph sparsity and structure, discrete diffusion models (Vignac et al., 2023; Kong et al., 2023) have been developed. However, a key challenge for these models is the slow sampling process due to the long reverse diffusion chain. To mitigate this limitation, several efficient diffusion techniques have been proposed, including EDGE (Chen et al., 2023), HiGen (Karami, 2024), ESGG (Bergmeister et al., 2024), and Pard (Zhao et al., 2024).
@@
Random walks have been widely used in graph learning due to their strong expressive power. GCKN (Chen et al., 2020) and RWGNN (Nikolentzos & Vazirgiannis, 2020) utilize path and walk kernels to learn graph representations. Several recent works (Ivanov & Burnaev, 2018; Wang et al., 2021; Yin et al., 2022) explicitly integrate random walk sequences with positional encodings, inspiring subsequent methods such as CRaWL (Tönshoff et al., 2023), NeuralWalker (Chen et al., 2024) and RWNN (Kim et al., 2024), which further enhance graph representation learning via random walk sequence modeling. GraphGPT (Zhao et al., 2023) leverages Eulerian paths to improve graph property prediction. Our work uniquely explores random sequence representations of graphs focusing on graph generation, introducing a novel perspective on combining random walks and language modeling for scalable graph generative modeling.
@@ 4 Experiments
Table 1: Comparison of AutoGraph to SOTA methods on Planar
Planar Graphs
$n_{\mathrm{graphs}}=128$, $|V|=64$
Model Deg. Clus. Orbit Spec. Ratio VUN
Training set 0.0002 0.0310 0.0005 0.0038 1.0
GraphRNN 0.0049 0.2779 1.2543 0.0459 638.5 0.0
GRAN 0.0007 0.0426 0.0009 0.0075 2.1 0.0
SPECTRE 0.0005 0.0785 0.0012 0.0112 2.6 25.0
EDGE 0.0761 0.3229 0.7737 0.0957 490.9 0.0
DiGress 0.0007 0.0780 0.0079 0.0098 6.1 77.5
ESGG 0.0005 0.0626 0.0017 0.0075 2.5 95.0
AutoGraph 0.0004 0.0605 0.0003 0.0064 1.5 87.5
@@
We employ the LLaMA model with 12 layers and a hidden dimension of 768 as our sequence model backbone across all experiments, aligning with the architecture of GPT-2’s smallest variant (Radford et al., 2019). Although prior works have used smaller models, we argue that our approach still demonstrates better scalability and faster training and inference speeds compared to diffusion models. For inference, we adopt the commonly used top-k sampling strategy (Fan et al., 2018). Our implementation leverages the Hugging Face framework (Jain, 2022), providing users with a flexible interface to experiment with SOTA language models for graph generation.
@@
For fair comparison, we align our evaluation methodology with established practices from prior works (You et al., 2018; Martinkus et al., 2022; Vignac et al., 2023). Our evaluation compares generated samples against the test set using maximum mean discrepancy (MMD) (Gretton et al., 2012), computed across multiple graph descriptors: node degree distributions (Deg.), clustering coefficients (Clus.), orbit count statistics (Orbit), and eigenvalue spectra (Spec.). As a reference, we also compute these metrics on the training set and report the mean ratio across all properties (Ratio).
For synthetic datasets, we additionally assess model performance using the VUN metric, the proportion of generated graphs that are valid, unique, and novel. Our efficiency analysis includes two measurements: inference speed, calculated as the per-graph generation time when producing 1024 graphs, and training efficiency, measured as the time required to achieve a VUN score of 75.0 for the Planar dataset and 60.0 for the SBM dataset. All efficiency measurements are performed on one NVIDIA H100 GPU.
For molecular generation datasets, we strictly follow the evaluation metrics used in DiGress (Vignac et al., 2023). More details about evaluation are provided in Appendix E.2.
@@ 4.1 Comparison to State-of-the-Art Methods
We evaluate the performance of AutoGraph compared to other SOTA graph generative models using the standard setting without pre-training.
Table 2: Comparison of AutoGraph to SOTA methods on SBM
Stochastic Block Models
$n_{\mathrm{graphs}}=128$, $|V|_{\max}=187$, $|V|_{\operatorname*{avg}}\approx 104$
Model Deg. Clus. Orbit Spec. Ratio VUN
Training set 0.0008 0.0332 0.0255 0.0027 1.0
GraphRNN 0.0055 0.0584 0.0785 0.0065 3.5 5.0
GRAN 0.0113 0.0553 0.0540 0.0054 5.0 25.0
SPECTRE 0.0015 0.0521 0.0412 0.0056 1.8 52.5
EDGE 0.0279 0.1113 0.0854 0.0251 12.7 0.0
DiGress 0.0018 0.0485 0.0415 0.0045 1.8 60.0
ESGG 0.0119 0.0517 0.0669 0.0067 5.4 45.0
AutoGraph 0.0077 0.0519 0.0439 0.0040 3.4 92.5
4.1.1 Small Synthetic Graph Generation
We first evaluate our method on the small synthetic graph datasets introduced by Martinkus et al. (2022), including the Planar and SBM datasets. We compare the performance of AutoGraph against GraphRNN (You et al., 2018), GRAN (Liao et al., 2019), SPECTRE (Martinkus et al., 2022), EDGE (Chen et al., 2023), DiGress (Vignac et al., 2023), ESGG (Bergmeister et al., 2024). As shown in Tables 1 and 2, AutoGraph achieves the best and second-best MMD ratios on average while ranking second-best and best in terms of VUN scores for the Planar and SBM datasets, respectively. Notably, all previous methods exhibit limited structural validity on the SBM dataset, with the best VUN scores reaching only 60.0.
Additionally, we assess the training and inference times of AutoGraph against representative models, including DiGress, GRAN, and ESGG. As presented in Table 3, AutoGraph is approximately 3 times faster during training and 100 times faster during inference compared to diffusion-based models. This substantial speedup over diffusion-based models is even more pronounced than that observed in other data modalities such as images (Tian et al., 2024).
@@ 4.1.1 Small Synthetic Graph Generation
Planar Training 25.9h OOT 7.4h 6.2h (4.2$\times$)
Inference 2.84s 0.03s 4.60s 0.01s (284$\times$)
SBM Training 47.7h OOT OOT 13.8h (3.5$\times$)
Inference 13.05s 0.13s 30.0s 0.14s (93$\times$)
Table 4: Comparison of AutoGraph to SOTA methods on the Proteins and Point Clouds datasets. OOM indices out of memory.
@@
GraphRNN 0.0040 0.1475 0.5851 0.0152 62.1 OOM OOM OOM OOM OOM
GRAN 0.0479 0.1234 0.3458 0.0125 77.7 0.0201 0.4330 0.2625 0.0051 19.1
SPECTRE 0.0056 0.0843 0.0267 0.0052 12.5 OOM OOM OOM OOM OOM
EDGE 0.1863 0.3406 0.6786 0.1075 274.5 0.4441 0.3298 1.0730 0.4006 104.7
DiGress 0.0041 0.0489 0.1286 0.0018 16.2 OOM OOM OOM OOM OOM
ESGG 0.0030 0.0309 0.0047 0.0013 4.7 0.0139 0.5775 0.0780 0.0055 6.8
@@
4.1.2 Large Graph Generation
To understand the scalability of AutoGraph, we evaluate its performance on the Proteins and Point Clouds datasets used by Liao et al. (2019). The results, shown in Table 4, demonstrate that even when using a context window shorter than the longest sequence sampled from the dataset, AutoGraph achieves MMD ratios comparable to those observed on the Planar and SBM datasets. Furthermore, AutoGraph outperforms all existing methods in terms of MMD ratio, achieving a twofold or more improvement over the previous best model, ESGG. More significantly, while ESGG was specifically designed for generating unattributed graphs, AutoGraph demonstrates versatility by being applicable to both unattributed and attributed graphs.
Table 5: Comparison of AutoGraph to SOTA methods on QM9
QM9 with hydrogen atoms
[…]
DiGress 95.4 97.6 33.4 98.1 79.8
AutoGraph 97.7 96.7 45.5 98.6 87.3
4.1.3 Molecular Graph Generation
We demonstrate the applicability of our method to generating real-world attributed graphs, such as molecular structures. We evaluate AutoGraph on the same datasets used by DiGress (Vignac et al., 2023), including QM9 (all atoms) (Wu et al., 2018), MOSES (Polykovskiy et al., 2020), and GuacaMol (Brown et al., 2019). Following the data splits and experimental setup from DiGress, we benchmark AutoGraph against a variety of SOTA models, including DiGress, VAE on SMILES (Polykovskiy et al., 2020), JT-VAE (Jin et al., 2018), GraphINVENT (Mercado et al., 2021), NAGVAE (Kwon et al., 2020), LSTM and MCTS (Brown et al., 2019). On the QM9 dataset (Table 5), AutoGraph outperforms DiGress across all metrics except uniqueness, showing its superiority for attributed graphs.
For the more challenging MOSES and GuacaMol datasets, AutoGraph also demonstrates superior performance, achieving higher validity and improved distributional alignment as measured by metrics like FCD, as shown in Tables 6 and 7. Notably, to our best knowledge, AutoGraph is the first autoregressive model for graphs to surpass diffusion-based approaches on these datasets. It is worth mentioning that all metrics were computed using SMILES representations rather than molecular graphs. Due to the non-reversible nature of converting SMILES to graphs and back, where approximately 20% of molecules cannot be mapped back to their original SMILES (Vignac et al., 2023), some discrepancies are introduced when calculating these metrics. Despite these challenges, AutoGraph achieves validity and FCD scores comparable to SMILES-based methods.
Furthermore, AutoGraph demonstrates remarkable efficiency, with training times of less than one day on both datasets, compared to up to one week for DiGress (Vignac et al., 2023). This substantial reduction in training time underscores AutoGraph’s practical advantages in large-scale and high throughput molecular graph generation tasks.
Table 6: Comparison of AutoGraph to SOTA methods on MOSES
@@ 4.1.3 Molecular Graph Generation
VAE SMILES 97.7 99.8 69.5 99.7 0.57 0.58
JT-VAE Fragments 100 100 99.9 97.8 1.00 0.53
GraphINVENT Graph 96.4 99.8 95.0 1.22 0.54
DiGress Graph 85.7 100 95.0 97.1 1.19 0.52
AutoGraph Graph 87.4 100 85.9 98.6 0.91 0.55
Table 7: Comparison of AutoGraph to SOTA methods on GuacaMol
@@
LSTM SMILES 95.9 100 91.2 99.1 91.3
NAGVAE Graph 92.7 95.5 100 38.4 0.9
MCTS Graph 100 100 99.4 52.2 1.5
DiGress Graph 85.2 100 99.9 92.9 68.0
AutoGraph Graph 91.6 100 97.7 97.5 79.2

### v2

@@ 2.3 Sampling Algorithm for SENT
Thm. 2.15 offers a simple strategy to sample a causal and Hamiltonian SENT: one needs to traverse the graph and choose the neighborhood set as all neighbors of the current node that have been visited. The traversing strategy could be achieved through a random path sampling or a depth-first search. In Algorithm 1, we provide a sampling strategy based on random path sampling with breaks.
Complexity analysis.
The length of a SENT, including the sizes of neighborhood sets (in other words, tokenized SENT defined in Section 2.4), is bounded by the number of nodes and edges, as each node and edge can be visited exactly once. Therefore, both the time and space complexity of sampling a SENT from graph $G$ are $\mathcal{O}(m)$ where $m$ is the number of edges.
2.4 Tokenization of SENT
@@
[…]
@@
[…]
$\texttt{Token}(s):=\texttt{Token}(s_{1})\mathbin{\|}[~\textbf{/}~]\mathbin{\|}\cdots\mathbin{\|}[~\textbf{/}~]\mathbin{\|}\texttt{Token}(s_{k}),\text{ where }\texttt{Token}(s_{i}):=\mathbin{\|}_{w\in s_{i}}\texttt{Token}(w),$
@@ 2.5 Extension to Attributed Graphs
Our method can be easily extended to graphs with categorical (or discretized) attributes by inserting node and edge attributes in an interleaved fashion into the tokenized SENT sequence. Specifically, let $L_{\mathrm{node}}(v)$ and $L_{\mathrm{edge}}(u,v)$ be the attributes of a node $v$ and an edge $(u,v)$ respectively. Using the same notation as above, we define for any $s_{i}:=(w_{1},\dots,w_{q})\in{\mathcal{T}^{\mathcal{N}}}$ with $w_{i}=(v_{i},\cdot)$:
$\texttt{Token}(s_{i})$ $=\texttt{Token}(w_{1})\mathbin{\|}[L_{\mathrm{edge}}(v_{1},v_{2})]\mathbin{\|}\texttt{Token}(w_{2})\mathbin{\|}\dots\mathbin{\|}\texttt{Token}(w_{q}),$
$\texttt{Token}(w)$ $=[v,L_{\mathrm{node}}(v),\bm{<},L_{\mathrm{edge}}(v,u_{1}),u_{1},\dots,L_{\mathrm{edge}}(v,u_{p}),u_{p},\bm{>}].$
@@ 2.6 Autoregressive Modeling of Tokenized SENTs
where the conditional probability $p_{\theta}$ is modeled using a neural network with parameters $\theta$. The architecture of the neural network can be any state-of-the-art sequence model.
@@ 3 Related Work
Autoregressive models generate graphs by sequentially adding nodes and edges. GraphRNN [66] pioneered this approach by framing graph generation as a sequence prediction task, demonstrating the capacity of recurrent neural networks (RNNs) [13] to capture complex structures. DeepGMG [38] introduced a probabilistic policy framework for conditional generation, while GRAN [39] and BiGG [14] enhanced efficiency and scalability by generating multiple nodes and edges in parallel.Recent research has focused on optimizing the generation order. Chen et al. [10] highlighted that the ordering of node and edge additions impacts graph quality, and GraphARM [36] applied reinforcement learning to dynamically refine this order. Goyal et al. [20] incorporated logical constraints to improve domain-specific generation, and Bacciu et al. [1] proposed Bayesian reasoning to better capture graph dependencies. BwR [16] and GEEL [29] investigated node ordering based on optimized bandwidth.These models, while efficient on synthetic datasets, do not explicitly represent graphs as token sequences, preventing direct application of LLM techniques. More significantly, their sequences are not guaranteed to be subgraph-induced (see Thm. 2.14). In contrast, our approach enables substructured-conditioned generation analogous to prompt-based generation in language models, establishing a more fundamental connection between graph and language modeling paradigms.
@@
Other graph generative models include variational, GAN-based, flow-based, and diffusion-based approaches. GraphVAEs [35, 53] employ variational autoencoders to learn latent representations, effectively generating small graphs but struggling with more complex structures. GAN-based models, such as NetGAN [4] and SPECTRE [41], generate graphs by modeling graph descriptors like random walks and spectral features. Flow-based methods such as [52] have shown the ability to generate small molecular graphs.Diffusion-based models iteratively refine noise into structured graphs through reverse diffusion steps. Continuous diffusion models [46, 31] adapt denoising diffusion probabilistic models for graph generation. To leverage graph sparsity and structure, discrete diffusion models [61, 36, 32, 64] have been developed. However, a key challenge for these models is the slow sampling process due to the long reverse diffusion chain. To mitigate this limitation, several efficient diffusion techniques have been proposed, including EDGE [11], HiGen [33], ESGG [2], and Pard [68].
@@
Random walks have been widely used in graph learning due to their strong expressive power. GCKN [8] and RWGNN [45] utilize path and walk kernels to learn graph representations. Several recent works [27, 62, 65] explicitly integrate random walk sequences with positional encodings, inspiring subsequent methods such as CRaWL [57], NeuralWalker [9] and RWNN [34], which further enhance graph representation learning via random walk sequence modeling. GraphGPT [69] leverages Eulerian paths to improve graph property prediction. Our work explores sequence representations of graphs for graph generation, introducing a novel perspective on combining random walks and language modeling for scalable graph generation.
@@ 4 Experiments
Table 1: Benchmarking AutoGraph on Planar and SBM
Planar Graphs Stochastic Block Models
$n_{\mathrm{graphs}}=128$, $|V|=64$ $n_{\mathrm{graphs}}=128$, $|V|_{\max}=187$, $|V|_{\operatorname*{avg}}\approx 104$
Model Deg. Clus. Orbit Spec. Ratio VUN Deg. Clus. Orbit Spec. Ratio VUN
Training set 0.0002 0.0310 0.0005 0.0038 1.0 0.0008 0.0332 0.0255 0.0027 1.0
GraphRNN [66] 0.0049 0.2779 1.2543 0.0459 638.5 0.0 0.0055 0.0584 0.0785 0.0065 3.5 5.0
GRAN [39] 0.0007 0.0426 0.0009 0.0075 2.1 0.0 0.0113 0.0553 0.0540 0.0054 5.0 25.0
SPECTRE [41] 0.0005 0.0785 0.0012 0.0112 2.6 25.0 0.0015 0.0521 0.0412 0.0056 1.8 52.5
EDGE [1] 0.0761 0.3229 0.7737 0.0957 490.9 0.0 0.0279 0.1113 0.0854 0.0251 12.7 0.0
GraphGen [20] 0.0328 0.2106 0.4236 0.0430 257.3 7.5 0.0550 0.0623 0.1189 0.0182 20.5 5.0
BiGG [14] 0.0007 0.0570 0.0367 0.0105 20.4 5.0 0.0012 0.0604 0.0667 0.0059 2.0 10.0
DiGress [61] 0.0007 0.0780 0.0079 0.0098 6.1 77.5 0.0018 0.0485 0.0415 0.0045 1.8 60.0
GruM [32] 0.0005 0.0353 0.0009 0.0062 1.8 90.0 0.0007 0.0492 0.0448 0.0050 1.5 85.0
GEEL [29] 0.0039 0.0013 0.0062 0.0234 9.5 0.0 0.0106 0.0616 0.0023 0.0381 7.3 5.0
ESGG [2] 0.0005 0.0626 0.0017 0.0075 2.5 95.0 0.0119 0.0517 0.0669 0.0067 5.4 45.0
AutoGraph 0.0004 0.0605 0.0003 0.0064 1.5 87.5 0.0077 0.0519 0.0439 0.0040 3.4 92.5
@@
We employ the LLaMA model with 12 layers and a hidden dimension of 768 as our sequence model backbone across all experiments, aligning with the architecture of GPT-2’s smallest variant [49]. Although prior works have used smaller models, we argue that our approach still demonstrates better scalability and faster training and inference speeds compared to diffusion models. For inference, we adopt the commonly used top-k sampling strategy [19]. Our implementation leverages the Hugging Face framework [28], providing users with a flexible interface to experiment with SOTA language models for graph generation.
@@
For fair comparison, we align our evaluation methodology with established practices from prior works [66, 41, 61]. Our evaluation compares generated samples against the test set using maximum mean discrepancy (MMD) [21], computed across multiple graph descriptors: node degree distributions (Deg.), clustering coefficients (Clus.), orbit count statistics (Orbit), and eigenvalue spectra (Spec.).As a reference, we also compute MMDs between the training and test sets and report the average ratio between generated and training MMDs (Ratio) following Bergmeister et al. [2].For synthetic datasets, we additionally assess model performance using the VUN metric, namely the proportion of generated graphs that are simultaneously valid, unique, and novel compared to the training graphs. Our efficiency analysis includes two measurements: inference speed, calculated as the per-graph generation time when producing 1024 graphs, and training speed, measured as the time required to achieve a VUN score of 75.0 for the Planar dataset and 60.0 for the SBM dataset. All efficiency measurements are performed on a single NVIDIA H100 GPU.For molecular generation datasets, we strictly follow the evaluation metrics used in DiGress [61] and use the evaluation tools from the official codebase [48, 6]. More details are provided in App. E.2.
@@ 4.1 Comparison to State-of-the-Art Methods
We evaluate the performance of AutoGraph compared to other SOTA graph generative models using the standard setting without pre-training.The comparison partners include GraphRNN [66], GRAN [39], SPECTRE [41], EDGE [11], GraphGen [20], BiGG [14], DiGress [61], GruM [32], GEEL [29], and ESGG [2].
Small synthetic graph generation.
@@
[…]
We first evaluate our method on the small synthetic graph datasets introduced by Martinkus et al. [41], including the Planar and SBM datasets.As shown in Table 1, AutoGraph demonstrates competitive MMDs while ranking second-best and best in terms of VUN scores on the Planar and SBM datasets, respectively. Importantly, only GruM and AutoGraph exhibit strong structural validity (VUN $\geq 80.0$) on the SBM dataset. Previous state-of-the-art autoregressive models, particularly GEEL, completely fail on both datasets in terms of VUN scores, as they largely memorize some subset of the training data. It is worth noting that the relatively low MMD ratio of AutoGraph is expected, as we selected the best model based on the VUN score. More results with error bars are provided in App. F.1.Additionally, we assess the training and inference times of AutoGraph against representative models, including DiGress, GRAN, and ESGG. As presented in Table 2, AutoGraph is approximately 3 times faster during training and 100 times faster during inference compared to diffusion-based models. This substantial speedup over diffusion-based models is even more pronounced than that observed in other data modalities such as images [56].
Table 3: Benchmarking AutoGraph on Proteins and Point Clouds. OOM indicates out of memory. Note that for GEEL [29], we fail to reproduce their experiments on the Point Clouds dataset using their official codebase.
@@
GraphRNN [66] 0.0040 0.1475 0.5851 0.0152 62.1 OOM OOM OOM OOM OOM
GRAN [39] 0.0479 0.1234 0.3458 0.0125 77.7 0.0201 0.4330 0.2625 0.0051 19.1
SPECTRE [41] 0.0056 0.0843 0.0267 0.0052 12.5 OOM OOM OOM OOM OOM
EDGE [1] 0.1863 0.3406 0.6786 0.1075 274.5 0.4441 0.3298 1.0730 0.4006 104.7
GraphGen [20] 0.0159 0.1677 0.3789 0.0181 58.1 OOM OOM OOM OOM OOM
BiGG [14] 0.0070 0.1150 0.4696 0.0067 50.1 0.0994 0.6035 0.3633 0.1589 38.2
DiGress [61] 0.0041 0.0489 0.1286 0.0018 16.2 OOM OOM OOM OOM OOM
GruM [32] 0.0019 0.0660 0.0345 0.0030 8.2 OOM OOM OOM OOM OOM
GEEL [29] 0.2110 0.3753 0.1768 0.1689 287.9
ESGG [2] 0.0030 0.0309 0.0047 0.0013 4.7 0.0139 0.5775 0.0780 0.0055 6.8
@@
Large graph generation.
To understand the scalability of AutoGraph, we evaluate its performance on the Proteins and Point Clouds datasets used by Liao et al. [39]. The results, shown in Table 3, demonstrate that even when using a context window shorter than the longest sequence sampled from the dataset, AutoGraph achieves MMD ratios comparable to those observed on the Planar and SBM datasets. Furthermore, AutoGraph outperforms all existing methods in terms of MMD ratio, achieving a twofold or more improvement over the previous best model, ESGG. More significantly, while ESGG was specifically designed for generating unattributed graphs, AutoGraph demonstrates versatility by being applicable to both unattributed and attributed graphs. Finally, previous state-of-the-art autoregressive models, such as GEEL, again fail to achieve competitive performance on these datasets.
Molecular graph generation.
We demonstrate the applicability of our method to generating real-world attributed graphs, such as molecular structures. We evaluate AutoGraph on the same datasets used by DiGress [61], including QM9 (all atoms) [63], MOSES [48], and GuacaMol [6]. Following the data splits and experimental setup from DiGress, we benchmark AutoGraph against a variety of SOTA models, including DiGress, VAE on SMILES [48], JT-VAE [30], GraphINVENT [43], NAGVAE [37], LSTM and MCTS [6]. On the QM9 dataset (Table 12), AutoGraph outperforms DiGress across all metrics except uniqueness, showing its superiority for attributed graphs.For the more challenging MOSES and GuacaMol datasets, AutoGraph also demonstrates superior performance, achieving higher validity and improved distributional alignment as measured by metrics like FCD, as shown in Table 4. Notably, to our best knowledge, AutoGraph is the first autoregressive model for graphs to surpass diffusion-based approaches on these datasets. It is worth mentioning that all metrics were computed using SMILES representations rather than molecular graphs. Due to the non-reversible nature of converting SMILES to graphs and back, where approximately 20% of molecules cannot be mapped back to their original SMILES [61], some discrepancies are introduced when calculating these metrics. Despite these challenges, AutoGraph achieves validity and FCD scores comparable to SMILES-based methods.Furthermore, AutoGraph demonstrates remarkable efficiency, with training times of less than one day on both datasets, compared to up to one week for DiGress [61]. This substantial reduction underscores AutoGraph’s practical advantages in large-scale molecular graph generation tasks.
Table 4: Benchmarking AutoGraph on the molecular generation datasets, more results in App. F.2. AutoGraph was first pretrained on the PubChem-10M dataset [12].
@@
VAE - SMILES 97.7 99.8 69.5 99.7 0.57 0.58
JT-VAE - Fragments 100 100 99.9 97.8 1.00 0.53
GraphINVENT - Graph 96.4 99.8 95.0 1.22 0.54
DiGress - Graph 85.7 100 95.0 97.1 1.19 0.52
AutoGraph - Graph 87.4 100 85.9 98.6 0.91 0.55
@@
LSTM - SMILES 95.9 100 91.2 99.1 91.3
NAGVAE - Graph 92.7 95.5 100 38.4 0.9
MCTS - Graph 100 100 99.4 52.2 1.5
DiGress - Graph 85.2 100 99.9 92.9 68.0
AutoGraph - Graph 91.6 100 97.7 97.5 79.2
AutoGraph - Graph 95.9 100 95.5 98.1 91.4